If the number of people classified as unemployed is 20,000 and the number of people classified as employed is 230,000, what is the unemployment rate?


1
Expert's answer
2021-10-27T08:30:14-0400

Solution:

Unemployment rate = "\frac{Unemployed\;people}{Labor\;force}\times 100"

Labor force = No. of employed persons + No. of unemployed persons

Unemployed people = 20,000

Employed people = 230,000

Labor force = 20,000 + 230,000 = 250,000


Unemployment rate = "\frac{20,000}{250,000} \times 100 = 8\%"